# IDEA: Enhancing the Rule Learning Ability of Large Language Model Agent through Induction, Deduction, and Abduction

### v5

@@ 5.3 Main Results
For the Oracle-rule agent, in the Escape Room puzzles, agents achieve up to an 89% success rate by simply following the provided rule. However, in other puzzles, merely knowing the rule is not sufficient for success; applying the rule to solve the puzzle is more challenging. The Baseline agent is not provided with the underlying rules and solely relies on historical observations to make interactive actions. Across models, the success rates drop by about half compared to the Oracle-rule agent. This significant decrease highlights the challenge of rule learning and indicates that current LLM agents struggle to learn rules in unfamiliar environments without explicit guidance.
IDEA significantly boosts success rates. Our proposed IDEA framework leads to approximate 10% increases in success rates for Llama3-70B, GPT-3.5-Turbo, and GPT-4o compared to the Baseline agent. This improvement demonstrates that incorporating a reasoning loop of abduction, deduction, and induction substantially enhances the LLM rule-learning performance in unfamiliar environments. IDEA enables the LLMs to generate hypotheses, plan actions, and refine their understanding based on new observations, which is crucial for rule learning. However, smaller models like Llama3-8B and Gemma-7B do not perform better when applying IDEA. This is likely due to the increased complexity and context management demands introduced by the IDEA framework, which pose challenges for small models with limited capacity.
@@ 5.4 Analysis
IDEA boosts puzzle-solving speed. Figure 3 illustrates the cumulative number of puzzles solved at each interaction step for the Baseline agent, the IDEA agent, and human participants. The slopes of the lines represent the rate at which puzzles are solved per step. Compared to the Baseline agent, the IDEA agent exhibits a steeper slope, indicating that the integration of abductive, deductive, and inductive reasoning enhances the agent’s efficiency in exploring the environment and learning the underlying rules, especially during the early stages.
Figure 3: Comparison of the cumulative number of puzzles solved at each interaction step. The IDEA agent significantly increases the solving speed compared to the Baseline agent.
@@
6 Human evaluation
@@
We employ three computer science graduate students (our co-authors) to evaluate all the hypotheses and plans generated by both the IDEA agent and human participants during the abduction, deduction, and induction stages within a randomly selected 50% subsample of the puzzles.
Abduction stage. In this stage, agents formulate an initial hypothesis based on initial observations. Since some puzzles are simple, agents can potentially guess the ground truth rule at this stage. Figure 4(a) shows that LLMs such as GPT-4o have approximately a 30% success rate in correctly identifying the rule during abduction. Notably, humans exhibit lower success rate in this stage. We believe this is because humans often lack the patience to thoroughly review all textual information initially and may overlook certain details. In contrast, LLMs meticulously process every word of the prompt and leverage their comprehensive pre-training data to generate plausible hypotheses. Additionally, humans tend not to formalize hypotheses when operating under uncertainty.
@@
Deduction stage. After establishing an initial hypothesis—or each time the agent refines a hypothesis—the agent derives a plan to either validate this hypothesis or use it to attempt the puzzle. As shown in Figure 4(b), humans generally outperform LLMs in creating high-quality plans.
Induction stage. Figure 4(c) illustrates the proportion of inductions where the refined hypothesis improved over the previous one, which we define as the effective induction rate. The induction stage is crucial in rule-learning since humans typically develop high-quality hypotheses through induction after experiments. However, LLM agents perform poorly at this stage. Humans significantly outperform LLMs, with 40% of their refined hypotheses being better than the previous ones. Surprisingly, GPT-4o performs worse than Llama3-70B and GPT-3.5-Turbo. We found that GPT-4o tends to hallucinate, especially in Escape Room puzzles, which negatively impacts its overall effective induction rate. In contrast, Llama3-70B almost never engages in induction within Reactor puzzles, as it fails to recognize when its hypothesized rules contradict its observations (see Appendix A.6.1).
@@ 7 Conclusion
In this work, we introduce RULEARN, a benchmark to evaluate LLM agents’ rule-learning abilities in interactive environments. We propose IDEA, an agent framework that mimics human reasoning through abduction, deduction, and induction. Comprehensive experiments involving five prominent LLMs and human participants reveal that while IDEA significantly improves the rule learning ability of LLM agents, there is still a large gap between LLM and humans particularly in refining hypotheses and adapting strategies. Despite these advancements from the IDEA framework, LLMs still face challenges in generating valid hypotheses and avoiding repetitive actions in complex scenarios. Our findings underscore the need for further development of LLMs that can emulate human cognitive processes more effectively in explorations of novel environments. RULEARN provides a foundational resource for future research aimed at closing these gaps.
@@ A.3.3 Pseudocode of interactive rule learning procedure
7:     $\mathbb{\tilde{M}}.\text{add}(\text{"You established a new {H} and {P}."})$" )
@@
27:              $\mathbb{\tilde{M}}.\text{add}(\text{"You established a new {H} and {P}."})$" )
@@ A.4 Computational Budget
For each setting (the Oracle agent, the Baseline, and the IDEA agent), we ran 300 different puzzles, with costs totaling \$188.10 for GPT-4o and \$37.62 for GPT-3.5. For open-source models, we utilized 8 RTX-A6000 GPUs. LLama3-70B required 5 days to complete its tasks, while Llama3-8b and Gemma-7b each took 2.5 days. We recruited 50 students as human participants, each working for one hour and receiving \$15 as compensation. In total, we spent over \$1,000 and utilized 80 GPU days for open-source models. The experimental cost is notably higher compared to traditional QA datasets due to each puzzle requiring over 20 steps and the rapid accumulation of context during solving.
@@ A.6 Human participants interface
Figure 18: An example of what human participants see during the experiment. Human participants view the same prompt as the LLM agents, except for additional overview information clarifying their responsibilities in the project. The name "{Nick_name}" is a pseudonym used to differentiate participants; we require participants to use only fictitious nicknames rather than their real names to ensure that the data collected is de-identified.
A.6.1 Hullicination examples

### v6

@@ 5.3 Main Results
For the Oracle-rule agent, in the Escape Room puzzles, agents achieve up to an 89% success rate by simply following the provided rule. However, in other puzzles, merely knowing the rule is not sufficient for success; applying the rule to solve the puzzle remains challenging. The Baseline agent is not provided with the underlying rules and solely relies on historical observations to make interactive actions. Across models, the success rates drop by about half compared to the Oracle-rule agent. This significant decrease highlights the challenge of rule learning and indicates that current LLM agents struggle to learn rules in unfamiliar environments without explicit guidance.
IDEA significantly boosts success rates. Our proposed IDEA framework leads to approximate 10% increases in success rates for Llama3-70B, GPT-3.5-Turbo, and GPT-4o compared to the Baseline agent. This improvement demonstrates that incorporating a reasoning loop of abduction, deduction, and induction substantially enhances the LLM rule-learning performance in unfamiliar environments. IDEA enables the LLMs to generate hypotheses, plan actions, and refine their understanding based on new observations, which is crucial for rule learning. However, smaller models like Llama3-8B and Gemma-7B do not perform better when applying IDEA. Small models inherently struggle with complex tasks like RULEARN —even when given the ground truth rule, their performance remains near 0—so no agent framework can significantly boost their performance.
@@ 5.4 Analysis
IDEA solves puzzles with fewer steps. Figure 4 illustrates the cumulative number of puzzles solved at each interaction step for the Baseline agent, the IDEA agent, and human participants. The slopes of the lines represent the rate at which puzzles are solved per step. Compared to the Baseline agent, the IDEA agent exhibits a steeper slope, indicating that the integration of abductive, deductive, and inductive reasoning enhances the agent’s efficiency in exploring the environment and learning the underlying rules at each interactive step, especially during the early stages.
Figure 4: Comparison of the cumulative number of puzzles solved at each interaction step. The IDEA agent significantly decrease number of steps needed to solve a puzzle compared to the Baseline agent.
@@
Table 4: Ablation on task-agnostic vs. task-specific prompts (GPT-4o, 50 % puzzle subset).
Setup All Types (%) Function Operator (%) Escape Room (%) Reactor (%)
Oracle-rule Agent 67.23 75.16 91.94 29.03
ReAct Agent (Baseline) 34.30 50.00 33.22 24.83
IDEA Agent (Task-Specific) 45.30 62.88 45.80 24.84
IDEA Agent (Task-Agnostic) 50.00 75.16 45.80 29.03
@@
IDEA is robust to different prompts. Recent work shows that large deep models often rely on shallow pattern matching and thus are sensitive to prompt changes and unable to generalize to new environments (Kang et al., 2024; Niu et al., 2024; Sun et al., 2024; Mirzadeh et al., 2024). For example, Mirzadeh et al. (2024) report that changing an irrelevant word in a math problem, for instance replacing “erasers” with “notebooks,” can greatly affect an LLM’s prediction.
As IDEA is expected to operate and learn new knowledge by interacting with unfamiliar environments, we conducted a supplementary experiment to test whether it remains effective under task-agnostic instructions. In our main experiments, we use task-specific instruction such as “Hypothesize the actual forms of each function.” In the supplementary experiment, we replaced that phrase with “Please consider the given observations and propose an initial hypothesis that explains them.” This prompt can apply equally to all rule-learning tasks (see Appendix A.5.4). We randomly sampled half of our puzzles and re-ran the evaluation with this new instruction. As shown in Table 4, IDEA still outperforms the baseline by a similar margin. Compared with other popular prompting method such as chain-of-thought (CoT) prompting (Wei et al., 2022), which typically requires carefully tailored prompts for each problem type, our single, high-level prompt works across all rule-learning tasks without further tuning. The agent first reviews its observations and then executes an abduction–deduction–induction loop, consistently delivering a performance boost.
6 Fine-grained Human Evaluation
@@
To compare LLM agent rule-learning with human processes, we conducted a fine-grained evaluation of rule-learning trajectories at every reasoning step. Computer science graduate students assessed the hypotheses and plans generated by both the IDEA agent and human participants during the abduction, deduction, and induction stages on a randomly selected 50% subsample of puzzles.
Abduction stage. In this stage, agents formulate an initial hypothesis based on early observations—sometimes even guessing the ground truth rule in simpler puzzles. Figure 5(a) indicates that LLMs such as GPT-4o correctly identify the rule during abduction about 30% of the time, while only around 10% of human participants generate a correct hypothesis under initial uncertainty. This discrepancy suggests that deviations in rule learning between LLMs and humans emerge as early as the abduction stage. LLMs tend to process every word of the prompt and produce a hypothesis even when unsure, whereas humans generally refrain from formalizing hypotheses under uncertainty.
@@
Deduction stage. After establishing an initial hypothesis—or each time the agent refines its hypothesis—the agent derives a plan to validate it or attempt the puzzle. As shown in Figure 5(b), humans generally outperform LLMs in creating high-quality plans. These superior plans enable humans to take a wider variety of actions, gathering more useful observations. According to Table LABEL:tab:repeat_count in Appendix A.2, humans ultimately collect 20% more diverse observations compared with LLMs.
Induction stage. Figure 5(c) shows the effective induction rate—the proportion of refined hypotheses that improve upon the previous version. Induction is crucial for developing high-quality hypotheses, and humans excel at this stage, with 40% of their refined hypotheses showing improvement. In contrast, LLMs struggle to converge on the correct rule, with effective induction rates below 20%. Moreover, they often fail to recognize conflicts between observations and hypotheses—for example, Llama3-70B rarely engages in induction within Reactor puzzles(see Appendix A.6.1).—resulting in redundant observations and fails the puzzle.
@@ 7 Conclusion
In this work, we introduce RULEARN, the first benchmark that (i) places LLM agents in fully interactive environments, (ii) gives them a fine-grained action space for discovering and applying rules, and (iii) employs manually designed rules that are unseen during pre-training. We propose IDEA, an agent framework that mimics human reasoning through abduction, deduction, and induction. Comprehensive experiments involving five prominent LLMs and human participants reveal that while IDEA significantly improves the rule learning ability of LLM agents, there is still a large gap between LLM and humans particularly in refining hypotheses and adapting strategies. Despite these advancements from the IDEA framework, LLMs still face challenges in generating valid hypotheses and avoiding repetitive actions in complex scenarios. Our findings underscore the need for further development of LLMs that can emulate human cognitive processes more effectively in explorations of novel environments. RULEARN provides a foundational resource for future research aimed at closing these gaps.
@@ 10 Acknowledgements
We thank the anonymous reviewers for their thoughtful comments. We are also grateful to Jia Li, Kevin Wang, and Kangshuo Li for their valuable feedback on puzzle design, and to Dr. Feng Chen for providing the computing resources necessary to complete the experiments.
@@ A.3.3 Pseudocode of interactive rule learning procedure
7:     $\mathbb{\tilde{M}}.\text{add}(\text{``You established a new {H} and {P}.''})$” )
@@
27:              $\mathbb{\tilde{M}}.\text{add}(\text{``You established a new {H} and {P}.''})$” )
@@ A.4 Computational Budget
For each setting (the Oracle agent, the Baseline, and the IDEA agent), we ran 300 different puzzles. On average, each API call’s input prompt is about 2,000 tokens, and the output is about 300 tokens. Completing each puzzle requires roughly 25 API calls (including both interaction and reasoning). Because each action changes the environment and updates the memory state, the input tokens cannot be cached, and each puzzle ends up using around 50,000 input tokens and 7,500 output tokens. With 300 puzzles in a single setting, that totals approximately 15,000,000 input tokens and 2,250,000 output tokens per setting. Since we test three settings (Oracle, Baseline, and IDEA), a single model uses about 45,000,000 input tokens and 6,750,000 output tokens across all puzzles and settings. For the open-source models, we used eight RTX-A6000 GPUs. LLaMA3-70B took about five days to complete its tasks, while LLaMA3-8B and Gemma-7B each required about 2.5 days. In total—across all three open-source models, three settings, and 300 puzzles—the open-source runs consumed roughly 80 GPU-days. The experimental cost is significantly higher compared to traditional QA datasets, primarily because each puzzle requires about 25 API calls in sequence and the context grows rapidly during solving. In the future, we plan to develop methods to reduce costs by organizing the input more efficiently, leveraging cached output capabilities, and incorporating a filtering mechanism to manage long-term memory.
@@ A.6 Human participants interface
Figure 18: An example of what human participants see during the experiment. Human participants view the same prompt as the LLM agents, except for additional overview information clarifying their responsibilities in the project. The name “{Nick_name}” is a pseudonym used to differentiate participants; we require participants to use only fictitious nicknames rather than their real names to ensure that the data collected is de-identified.
A.6.1 Hallucination examples